OpenStudy (anonymous):

Solving: 7x - 4 > 31 and 7x - 4 <= 59 7x > 35 and 7x <= 63 x > 5 and x <= 9 Is that right?

OpenStudy (anonymous):

yes, you shouldnt doubt yourself

OpenStudy (anonymous):

I can't help it, lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

hey for the one you did for me before: -3x + 4 <= x + 8 4 <= 3x + 8 -4 <= 3x -4/3 = x How would you describe the graph of the solution?

OpenStudy (anonymous):

where did you go? lol